# SuperPrimitive: Scene Reconstruction at a Primitive Level

### v1

@@ 4.2 Few-View Structure-from-Motion
Figure 4: 3-View SfM on ScanNet. We provide the visualisations of unprojected reference frame depth maps after predicted by our method for few-view SfM using one reference and $2$ supplementary views. Note that we used surface normal prediction network which was only pretrained on HyperSim [38] for this experiment.
@@ 4.3 Monocular Visual Odometry
We compare against other VO systems that do not have global bundle adjustment, following the protocol of [10] and evaluate on 8 sequences from the Frieburg 1 split. We use only RGB images as the input to our system and downsample them to $120\times 160$ for efficiency purposes. Despite its simplicity, our system outperforms all other methods in terms of Average Trajectory Error (ATE) [44]. Since our method is purely monocular, the estimated trajectory lacks global scale and we first use $\mathrm{Sim}\!\left(3\right)$ alignment to the ground truth scale, following standard practice [4]. As shown in the results in Tab. 2, despite our method simplicity our system outperforms all other methods in term of average ATE across all trajectories. Additionally, our VO is either the best or second best on five out of eight sequences. Besides quantitative evaluation, we also demonstrate reconstruction results in Fig. 5.
@@ 5 Conclusion
We presented a new representation, SuperPrimitives, which demonstrates how recent advances in building strong single image priors could be incorporated into pose and dense geometry estimation problems. We show that incorporating these priors alleviates the need for sophisticated hand-crafted heuristics and paves the way into monocular reconstruction with relative ease.
6 Acknowledgments

### v2

@@ 4.2 Few-View Structure-from-Motion
Figure 4: 3-View SfM on ScanNet. We provide the visualisations of unprojected reference frame depth maps predicted by our method for few-view SfM using one reference and $2$ supplementary views. Note that we used surface normal prediction network which was only pretrained on HyperSim [40] for this experiment.
@@ 4.3 Monocular Visual Odometry
We compare against other VO systems that do not have global bundle adjustment, following the protocol of [10] and evaluate on 8 sequences from the Frieburg 1 split. We use only RGB images as the input to our system and downsample them to $120\times 160$ for efficiency purposes. Since our method is purely monocular, the estimated trajectory lacks global scale and we first use $\mathrm{Sim}\!\left(3\right)$ alignment to the ground truth scale, following standard practice [4]. Tab. 2 shows that, despite the simplicity of our odometry system, it outperforms all other methods in terms of Average Trajectory Error (ATE) [46], averaged across all trajectories. Additionally, our VO is either the best or second best on five out of eight sequences. Besides quantitative evaluation, we also demonstrate reconstruction results in Fig. 5.
@@ 5 Conclusion
We presented a new representation, SuperPrimitive, which demonstrates how recent advances in building strong single image priors could be incorporated into pose and dense geometry estimation problems. We show that incorporating these priors alleviates the need for sophisticated hand-crafted heuristics and paves the way into monocular reconstruction with relative ease.
Acknowledgments.
@@
[…]
Supplementary Material
6 Ablation Study
6.1 Surface Normal vs Depth Prior
While a monocular depth prior has been widely used in 3D reconstruction [2, 8], in this work we use a surface normal prior instead. This choice is driven by the fact that surface normals have stronger generalisation abilities and capture better scene geometry. This observation is confirmed with an ablation study on the depth completion experiment described in  Fig. 3.
In this ablation experiment, we replaced our unscaled depth estimates $\mathfrak{D}$ (obtained by integrating surface normals) with the outputs of DPT-Hybrid [37], a state-of-the-art depth estimator. All depth scales are then optimised in the same way as in the original method.
Our approach marginally outperforms its DPT counterpart on the VOID depth completion benchmark on all metrics (Tab. 4). Qualitatively (Fig. 7), our method is better at preserving structural properties of the scene, such as walls.
6.2 Surface Normal Quality Impact
We investigate how the surface normal quality affects the performance of our method, especially for pose estimation. There are two ablation levels we performed in this case. The first one just assumes constant depth $z=const$ within each SuperPrimitive. This makes our method akin to Multiplane images (MPI) [44, 59]. To emulate planar but possibly slanted segments, we replaced surface normal vectors with their averaged value within each SuperPrimitive independently. Our full odometry system performs significantly better than its two ablated counterparts, see Tab. 3.
Method ATE
Constant Depth 0.257
Constant Normals 0.208
Ours 0.153
Table 3: Ablation study on TUM. We ablate our odometry by flattening out SuperPrimitives. See the main text for the details.
Method MAE RMSE iMAE iRMSE
Ours With Depth From DPT-Hybrid 128.03 213.32 57.86 91.82
Ours 109.0 204.15 47.32 83.40
Table 4: Ablation Study on VOID. The quality of the unscaled depth obtained from method via surface normal integration is quantitatively compared against direct monocular depth estimation with DPT-Hybrid.
Figure 7: Qualitative Ablation on VOID.
7 Implementation Details
Segments Post Processing.
Segments $\Omega_{i}$ extracted from the segmentation model may not be connected a priori. We perform a simple mask connectivity check and split a primitive into two in the case of detected mask discontinuities.
Depth scales.
Our depth scaling is implemented via storing a point $p_{i}$ within the segment $\Omega_{i}$. We represent depth scales $s_{i}$ as the log-depth value at $p_{i}$. This point $p_{i}$ is the same as the query point provided to the segmentation model as an input. This depth scale parametrisation allows converting partially available depth maps into a set of depth scaled SuperPrimitives.
7.1 Few-View Structure-From-Motion
We initialised depth scales uniformly to $1.0$ for each primitive. All supplementary poses are initialised at identity. We solve SfM in a coarse-to-fine fashion to ensure the segments would not stuck in a local minima.
A small penalty on depth scale was added with the weight $w=1\mathrm{e}{-5}$ to constraint segments that may not have photometric information from other views.
7.2 Depth Reinitialisation in MonoVO
Given a new keyframe $\mathrm{I}^{\text{next}}_{\text{kf}}$ with an estimated pose $\mathrm{T}^{\text{next}}_{\text{kf}}$, we scale depth for each new SuperPrimitve by using the geometry estimates of the previous keyframe $\mathrm{I}^{\text{prev}}_{\text{kf}}$.
More precisely, we transform the point cloud $\mathcal{G}_{\text{prev}}$ of the previous keyframe into the coordinate system of the new keyframe and then render a partial depth map. Then, the depth scales $s_{i}$ of the new keyframe are estimated as in the depth completion experiments in Fig. 3.
8 Experimental details
8.1 VOID Dataset
For depth completion evaluation on the VOID dataset, we follow the protocol of [55]. The ground truth depth is considered to be valid between $0.2$ and $5.0$ meters. The test set consists of $800$ images. The dataset also provides sparse depth measurements obtained by an external visual-inertial odometry. We choose the setting with least sparse depth measurements available, “150 points” — where the SLAM system was configured to estimate depth of around $150$ feature points (which constitutes $0.05\%$ of the full image size).
Metric Units Definition
MAE mm $\frac{1}{|\Omega|}\sum_{\bm{u}\in\Omega}|\hat{z}(\bm{u})-z_{\text{gt}}(\bm{u})|$
RMSE mm $\big{(}\frac{1}{|\Omega|}\sum_{\bm{u}\in\Omega}|\hat{z}(\bm{u})-z_{\text{gt}}(\bm{u})|^{2}\big{)}^{1/2}$
iMAE 1/km $\frac{1}{|\Omega|}\sum_{\bm{u}\in\Omega}|1/\hat{z}(\bm{u})-1/z_{\text{gt}}(\bm{u})|$
iRMSE 1/km $\big{(}\frac{1}{|\Omega|}\sum_{\bm{u}\in\Omega}|1/\hat{z}(\bm{u})-1/z_{\text{gt}}(\bm{u})|^{2}\big{)}^{1/2}$
Table 5: Error metrics. The definition of the error metrics used in depth quality valuation. Here, $\hat{z}$ and $z_{gt}$ are predicted and ground truth depth values respectively.
8.2 TUM RGB-D Dataset
For evaluation of the estimated trajectory we used ATE RMSE metric [46].
9 Hardware Details
All of our experiments were conducted in the following hardware setup: Intel Core i7 3.60GHz processor, 32 GB RAM, and NVIDIA GeForce RTX 4090 with 24GB VRAM. Our method is implemented in PyTorch [35] and CuPy [33] libraries.